OpenStudy (anonymous):

It is being given that 1 is one of the zeroes of the polynomial 7x-x^3-6.Find its other zeroes

OpenStudy (anonymous):

 \[1*-6=-6,3-2=1,3*-2=-6\] \[x ^{2}+3x-2x-6=0\] make factors and find the values of x,these are other zeroes.

OpenStudy (anonymous):

I didnt get :\

OpenStudy (anonymous):

\[x \left( x+3 \right)-2\left( x+3 \right)=0\] \[\left( x+3 \right)\left( x-2 \right)=0\] now can you solve?

OpenStudy (anonymous):

Okay ._. x+3=0 x=-3 Similarly, x-2=0 x=2 Is that right?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Yay :D

OpenStudy (anonymous):

Thank you :)